# Arbitrary Waveform Generator CW Continuous Wave

### v1

@@
Faruk Pasic1, Lukas Eller12, Stefan Schwarz1, Markus Rupp1 and Christoph F. Mecklenbräuker1 This work has been funded by the Christian Doppler Laboratory for Digital Twin assisted AI for sustainable Radio Access Networks, Institute of Telecommunications, TU Wien. The financial support by the Austrian Federal Ministry for Labour and Economy and the National Foundation for Research, Technology and Development and the Christian Doppler Research Association is gratefully acknowledged. 1Institute of Telecommunications, TU Wien, Vienna, Austria 2Christian Doppler Laboratory for Digital Twin assisted AI for sustainable Radio Access Networks faruk.pasic@tuwien.ac.at
@@ Abstract
Index Terms:
mmWave, MIMO, channel estimation, deep learning, CNN, UNet.
@@ I Introduction
[…]
@@ IV Simulation-based Comparison
MIMO Configuration 8$\times$8 8$\times$8
@@
where $L_{r}$ denotes the number of realizations and $l$ indicates a specific channel realization. The achievable SE, averaged over $N^{\left(\rm m\right)}$ subcarriers and expressed in bits///s///Hz, is defined by
@@
Results at K-factor of --20 dB: In this scenario, the in-band and out-of-band aided ML methods exhibit comparable performance, all outperforming the MRC method across the entire SNR range.
@@
SNR $\gamma^{\left(\rm m\right)}$ [dB] $\mathcal{U}\sim$ [--20, 10]
$K$-factor $K^{\left(\rm m\right)}$ [dB] $\mathcal{U}\sim$ [--20, 30]
AoD $\vartheta$, AoA $\varphi$ [deg] $\mathcal{U}\sim$ [--90, 90]

### v2

@@
Faruk Pasic1, Lukas Eller12, Stefan Schwarz1, Markus Rupp1 and Christoph F. Mecklenbräuker1 This work has been funded by the Christian Doppler Laboratory for Digital Twin assisted AI for sustainable Radio Access Networks, Institute of Telecommunications, TU Wien. The financial support by the Austrian Federal Ministry for Labour and Economy and the National Foundation for Research, Technology and Development and the Christian Doppler Research Association is gratefully acknowledged.
@@ I Introduction
Notation: The superscript $\left(\cdot\right)^{\left(\rm b\right)}$ represents frequency-band dependent values, where ${\rm b}\in\{{\rm s},{\rm m}\}$, with $\rm s$} referring to the sub-6 GHz and $\rm m$} to the mmWave band. Scalars are denoted by $x$, vectors by bold lowercase letters $\mathbf{x}$} and matrices by bold uppercase letters $\mathbf{X}$}. The $i$-th row of the matrix $\mathbf{X}$} is indicated by $\mathbf{X}_{i,:}$, and the $j$-th column by $\mathbf{X}_{:,j}$. The all-ones matrix is denoted by $\mathbf{1}$, with its dimensions specified in the subscript. The superscripts $\left(\cdot\right)^{\rm T}$ and $\left(\cdot\right)^{\rm H}$ represent transpose and Hermitian transpose, respectively. The Euclidean norm is represented by $\lVert\cdot\rVert$ and the Frobenius norm by $\lVert\cdot\rVert_{F}$.
@@ IV Simulation-based Comparison
MIMO Configuration $\times8$ $\times8$
@@
where $L_{r}$ denotes the number of realizations and $l$ indicates a specific channel realization. The achievable SE, averaged over $N^{\left(\rm m\right)}$ subcarriers and expressed in bits//s//Hz, is defined by
@@
Results at K-factor of -20 dB: In this scenario, the in-band and out-of-band aided ML methods exhibit comparable performance, all outperforming the MRC method across the entire SNR range.
@@
SNR $\gamma^{\left(\rm m\right)}$ [dB] $\mathcal{U}\sim$ [-20, 10]
$K$-factor $K^{\left(\rm m\right)}$ [dB] $\mathcal{U}\sim$ [-20, 30]
AoD $\vartheta$, AoA $\varphi$ [deg] $\mathcal{U}\sim$ [-90, 90]